OpenStudy (johan14th):

How do you solve for the derivative of (e^2x)/2x

OpenStudy (aravindg):

It is of the form f(x)/g(x) Use quotient rule.

OpenStudy (johan14th):

Ive set it up as \[\frac{ 2x(2xe ^{2x})- e ^{2x}(2) }{ 2x^2 }\]

OpenStudy (johan14th):

i have tried reducing it but its not any of the choices on my homework. they seem to factor a 2 or 2x out but i cant get the inside.

OpenStudy (caozeyuan):

downstair shoudl be 4x^2?

OpenStudy (caozeyuan):

I am rusty on calc though

OpenStudy (johan14th):

okay, i will try that

OpenStudy (johan14th):

hmm... I cant seem too get an any of the answers using \[4x^2\]either...

OpenStudy (caozeyuan):

what are the choices?

OpenStudy (johan14th):

a. \[\frac{ 2(2xe-2) }{ x }\]

OpenStudy (johan14th):

b. \[\frac{ 2e(2x-2) }{ 2x }\]

OpenStudy (caozeyuan):

what? e on its own, not e^x or something?

OpenStudy (johan14th):

ohh sorry i forgot a the ^2, it should be 2e^2

OpenStudy (johan14th):

c. \[\frac{ 4x(e^2-2) }{ 2x^2 }\]

OpenStudy (caozeyuan):

I'll go ask Wolfram

OpenStudy (johan14th):

Thanks

terenzreignz (terenzreignz):

你好 Nei Hou everyone XD @caozeyuan Giving up so soon? :>

OpenStudy (johan14th):

d. \[\frac{ 2x(e^2-2)}{ 2x^2 }\]

OpenStudy (caozeyuan):

\[\frac{ e^{2x}(2x-1) }{ 2x ^{2} }\]

OpenStudy (anonymous):

As it's already been mentioned, we use the quotient rule: \[\dfrac{d}{dx}(\dfrac{f(x)}{g(x)}) = \dfrac{f'(x)g(x)-f(x)g'(x)}{g(x)^2}\]Where \[f(x)=e^{2x}\]\[g(x)=2x\]So\[f'(x)=2e^{2x}\]\[g'(x)=2\]Giving us\[\dfrac{2e^{2x}\times2x-e^{2x}\times2}{4x^2} = \dfrac{4xe^{2x}-2e^{2x}}{4x^2} = \dfrac{(2x-1)e^{2x}}{2x^2}\]

OpenStudy (anonymous):

That last answer you posted is right @caozeyuan

OpenStudy (aravindg):

Its (2x)^2 not 2x^2

terenzreignz (terenzreignz):

A typo in the choices, then?

OpenStudy (johan14th):

Okay i see that i think the e^2x confused me a bit in my math. Thank you :) @tom982

OpenStudy (caozeyuan):

Yep, none of the choices make sense to me

OpenStudy (anonymous):

No problem Johan :)

OpenStudy (johan14th):

It could be that my teacher was excited for the thanksgiving break. Took a test on our last day some of the answers where repeated... so not a very reliable way to check my work.

OpenStudy (johan14th):

Thanks everyone, i will trudge through the rest of my homework now C:

OpenStudy (anonymous):

Hope it goes well. Drop me a message if you want me to have a look at something.